OpenStudy (anonymous):

Solve. Please show the algebraic inequality you used and show all of your work. Donna has the following scores on her accounting tests: 92, 83, 95, 89, 82. What does she have to score on her last test in order to end the term with a A average? An A for her university starts at 90.

OpenStudy (anonymous):

add all the scores together and add x. Divided that all by 6 and make it equal to 90. Solve for x.

OpenStudy (anonymous):

92+83+95+89+82+X/6 equal or greater than 90 x equal or greater than 99 to have an A

OpenStudy (anonymous):

Let x denote the score on the last test. We see that there are six tests total, so the average will be all of them added together, divided by 6: average = 92 + 83 + 95 +89 + 82 + x / 6 = 441 + x / 6 this has to be greater than 90: 441 + x / 6 > 90 we subtract 441 / 6 from both sides to get: x/6 > 90 - 441/6 = 90 - 73.5 = 16.5 We multiply both sides by 6 to get: x > 99 So she'll have to get 99% or higher.

OpenStudy (anonymous):

thats what i said x is equal or greater than 99

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

that's it